# Ringdown in Vaidya spacetimes: time-dependent frequencies, Penrose limit and time-domain analyses

### v2

@@ Acknowledgements
This work was supported by JSPS KAKENHI Grant Numbers JP25K07281 (C.Y.), JP24K07027 (C.Y.), JP22K03626 (M.K.), JP21H05182 (A.I.), JP21H05186 (A.I.), and JP25K07306 (A.I.).
@@ 1. Penrose limit geometries
First, let us briefly review the Penrose limit geometries [16] following Ref. [17]. The procedure to obtain the Penrose limit metric around a null geodesic $\gamma$ starts from finding the Penrose coordinates adapted to $\gamma$ in which the metric is described by
$\mathrm{d}s^{2}_{\gamma}=2\mathrm{d}U\mathrm{d}V+a(U,V,Y^{k})\mathrm{d}V^{2}+2b_{i}(U,V,Y^{k})\mathrm{d}Y^{i}\mathrm{d}V+g_{ij}(U,V,Y^{k})\mathrm{d}Y^{i}\mathrm{d}Y^{j},$ (134)
where $i$, $j$ and $k$ run over 1 to 2. There is a general way to find this Penrose coordinate based on the Hamilton-Jacobi function $S(x^{\mu})$ for null geodesic congruence involving $\gamma$, which will be briefly described in the next subsection. Readers are asked to refer to Ref. [17] for details. We will also describe the specific example of the construction of the Penrose coordinates for the unstable circular orbit of photons (UCOP) in a static spherically symmetric spacetime in  B. However, since we will take a shortcut to obtain the Penrose limit geometry in the main text, the derivation is not necessarily needed for understanding our results in the main text.
Once the metric in the Penrose coordinates is given, the Penrose limit is defined by the following procedure:
$\mathrm{d}\bar{s}^{2}=\lim_{\varepsilon\rightarrow 0}\left(\varepsilon^{-2}\left.\mathrm{d}s_{\gamma}^{2}\right|_{(U,V,Y^{k})\rightarrow(U,\varepsilon^{2}V,\varepsilon Y^{k})}\right).$ (135)
@@
$\mathrm{d}\bar{s}^{2}=2\mathrm{d}U\mathrm{d}V+\bar{g}_{ij}(U)\mathrm{d}Y^{i}\mathrm{d}Y^{j},$ (136)
where $\bar{g}_{ij}(U)=g_{ij}(U,0,0)=\lim_{\varepsilon\rightarrow 0}g_{ij}(U,\varepsilon^{2}V,\varepsilon Y^{k})$. This metric form is known as a plane wave metric in Rosen coordinates. It is worth noting that the finite region described by the cubic domain of the coordinates $(U,V,Y^{k})$ for the Penrose limit metric corresponds to the infinitely thin region along the geodesic $\gamma$ with a finite length of the affine parameter $U$. Therefore, taking the Penrose limit, we can extract the geometry in the vicinity of the null geodesic $\gamma$.
@@
$\mathrm{d}\bar{s}^{2}=2\mathrm{d}u\mathrm{d}v+A_{ab}(u)x^{a}x^{b}\mathrm{d}u^{2}+\delta_{ab}\mathrm{d}x^{a}\mathrm{d}x^{b},$ (137)
@@
$\bar{E}^{i}_{(b)}\partial_{u}\bar{E}^{(a)}_{i}=\bar{E}^{i}_{(a)}\partial_{u}\bar{E}^{(b)}_{i}$. (141)
Then the metric components $A_{ab}$ is given by
$A_{ab}=\delta_{ac}\ddot{\bar{E}}^{c}_{i}\bar{E}_{b}^{j}.$ (142)
@@
$g_{\mu\nu}E^{i}_{(a)}(\partial_{i})^{\nu}E^{j}_{(b)}(\partial_{j})^{\nu}=E_{j}^{(a)}E^{j}_{(b)}=\delta^{a}_{b},$ (143)
$(\partial_{U})^{\mu}\nabla_{\mu}(E^{i}_{(a)}(\partial_{i})^{\nu})=0.$ (144)
@@
The identification between $\bar{E}^{(a)}_{i}$ and $E^{(a)}_{i}$ enable us to make a covariant description of the Penrose limit. The covariant description can be made based on the following two facts:
$\bar{R}_{UiU}^{~~~~j}=\left.R_{UiU}^{~~~~j}\right|_{\gamma},$ (145)
$A_{ab}=-\bar{R}_{UiUj}\bar{E}^{i}_{(a)}\bar{E}^{j}_{(b)}=-\left.\left(R_{UiUj}E^{i}_{(a)}E^{j}_{(b)}\right)\right|_{\gamma},$ (146)
where $\bar{R}$ describes the Riemann tensor for $\bar{g}_{ij}$. We note that the first equation follows the fact that $R_{UiU}^{~~~~j}$ for the metric (134) does not depend on $a(U,V,Y^{k})$ nor $b_{i}(U,V,Y^{k})$. Since the Penrose limit metric in the Brinkmann coordinates is totally characterized by $A_{ab}$, we can obtain the metric form of Eq. (137) by calculating the Riemann tensor components and the parallelly propagated zweibein in the metric (134) without performing the explicit coordinate transformations.
@@ 2. Construction of the Penrose coordinates based on the Hamilton-Jacobi function
$g_{UU}$ $=$ $g_{\mu\nu}\frac{dx^{\mu}}{d\tau}\frac{dx^{\nu}}{d\tau}=0,$ (150)
$g_{UV}$ $=$ $g_{\mu\nu}\frac{\partial x^{\mu}}{\partial U}\frac{\partial x^{\mu}}{\partial V}=\partial_{\mu}S\frac{\partial x^{\mu}}{\partial V}=\frac{\partial S}{\partial V}=1,$ (151)
$g_{Ui}$ $=$ $g_{\mu\nu}\frac{\partial x^{\mu}}{\partial U}\frac{\partial x^{\mu}}{\partial Y^{i}}=\frac{\partial S}{\partial Y^{i}}=\frac{\partial V}{\partial Y^{i}}=0.$ (152)
@@ Appendix B Penrose limit for a circular photon orbit
$\mathcal{L}=\frac{1}{2}\left(-f\dot{t}^{2}+h\dot{r}^{2}+r^{2}\dot{\theta}^{2}\right).$ (153)
@@
$S(x^{\mu})=-Et+L\theta+\mathcal{\mathcal{}}R(r).$ (154)
In Ref. [17], the coordinate construction is given for the Schwarzschild metric with the initial hyper-surface being $r={\rm const.}$ surface. However, this surface $r={\rm const.}$ is tangent to the characteristic curve $r=r_{\rm c}$, and does not match the case we suppose. Although the procedure shown in Ref. [17] gives the same Penrose limit metric in Brinkmann coordinates as ours, we avoid this subtlety by taking a different initial hyper-surface. Specifically, we implicitly assume that the initial hypersurface is given by $F(\theta,r)=0$.
@@
$\mathrm{d}\theta$ $=$ $\dot{\theta}\mathrm{d}U+(\dot{\theta}-\beta)\mathrm{d}\tilde{r}$, (155)
$\mathrm{d}r$ $=$ $\dot{r}\mathrm{d}U+\dot{r}\mathrm{d}\tilde{r}$, (156)
@@
$\mathrm{d}S=\mathrm{d}V=-E\mathrm{d}t+L\mathrm{d}\theta+R^{\prime}(r)\mathrm{d}r$ (157)
$\Leftrightarrow$ $\mathrm{d}t=-\frac{1}{E}\mathrm{d}V+\frac{\beta^{2}\Delta^{2}}{E^{2}}\mathrm{d}U+\left[\frac{\Delta^{2}}{Efr}-\frac{L}{E}\left(\frac{L}{r^{2}}-\beta^{2}\right)\right]\mathrm{d}\tilde{r}$, (158)
@@
$\mathrm{d}\bar{s}^{2}$ $=$ $2\mathrm{d}U\mathrm{d}V+\frac{\beta^{2}\Delta^{2}}{E^{2}}\mathrm{d}\tilde{r}^{2}+r^{2}\sin^{2}\theta\mathrm{d}\tilde{\phi}^{2}$ (160)
@@
$\mathrm{d}\bar{s}^{2}=2\mathrm{d}U\mathrm{d}V+\frac{\beta^{2}\Delta^{2}(r(U))}{E^{2}}\mathrm{d}\tilde{r}^{2}+r^{2}(U)\sin^{2}\left(L\int r^{-2}(U)\mathrm{d}U\right)\mathrm{d}\tilde{\phi}^{2},$ (161)
@@
$E_{\mu}^{(\tilde{r})}=\frac{\beta\Delta(r(U))}{E}(\mathrm{d}\tilde{r})_{\mu}~~\&~~E_{\mu}^{(\tilde{\phi})}=r(U)\sin\left(L\int r^{-2}(U)\mathrm{d}U\right)(\mathrm{d}\tilde{\phi})_{\mu}.$ (162)
The metric components $A_{ab}$ can be easily calculated using Eq. (142). In the limit approaching a UCOP, we obtain
$A_{11}$ $=$ $\left.\frac{\ddot{\Delta}}{\Delta}\right|_{r=r_{\rm c}}=-V_{c}^{\prime\prime}=\alpha^{2},$ (163)
$A_{22}$ $=$ $\left.\frac{(r\sin\theta)\ddot{~}}{(r\sin\theta)}\right|_{r=r_{\rm c}}=-\left.\dot{\theta}^{2}\right|_{r=r_{\rm c}}=-\beta^{2}.$ (164)
@@
$t$ $\hat{=}$ $\frac{E}{f_{\rm c}}(u+\Lambda(V))-\frac{v}{E},$ (165)
$\theta$ $\hat{=}$ $\beta(u+\Lambda(V)),$ (166)
where we have used the symbol $\hat{=}$ to indicate the equality only on the UCOP. These relations correspond to Eqs. (19) and (22).

### v3

@@ Acknowledgements
This work was supported by JSPS KAKENHI Grant Numbers JP25K07281 (C.Y.), JP24K07027 (C.Y.), JP22K03626 (M.K.), JP21H05182 (A.I.), JP21H05186 (A.I.), and JP25K07306 (A.I.). R.O. would like to take this opportunity to thank the “THERS Make New Standards Program for the Next Generation Researchers” supported by JST SPRING, Grant Number JPMJSP2125.
@@ 1. Penrose limit geometries
First, let us briefly review the PL geometries [16] following Ref. [17]. The procedure to obtain the PL metric around a null geodesic $\gamma$ starts from finding the Penrose coordinates adapted to $\gamma$ in which the metric is described by
$\mathrm{d}s^{2}_{\gamma}=2\mathrm{d}U\mathrm{d}V+a(U,V,Y^{k})\mathrm{d}V^{2}+2b_{i}(U,V,Y^{k})\mathrm{d}Y^{i}\mathrm{d}V+g_{ij}(U,V,Y^{k})\mathrm{d}Y^{i}\mathrm{d}Y^{j},$ (A.1)
where $i$, $j$ and $k$ run over 1 to 2. There is a general way to find this Penrose coordinate based on the Hamilton-Jacobi function $S(x^{\mu})$ for null geodesic congruence involving $\gamma$, which will be briefly described in the next subsection. Readers are asked to refer to Ref. [17] for details. We will also describe the specific example of the construction of the Penrose coordinates for the unstable circular orbit of photons (UCOP) in a static spherically symmetric spacetime in  B. However, since we will take a shortcut to obtain the PL geometry in the main text, the derivation is not necessarily needed for understanding our results in the main text.
Once the metric in the Penrose coordinates is given, the PL is defined by the following procedure:
$\mathrm{d}\bar{s}^{2}=\lim_{\varepsilon\rightarrow 0}\left(\varepsilon^{-2}\left.\mathrm{d}s_{\gamma}^{2}\right|_{(U,V,Y^{k})\rightarrow(U,\varepsilon^{2}V,\varepsilon Y^{k})}\right).$ (A.2)
@@
$\mathrm{d}\bar{s}^{2}=2\mathrm{d}U\mathrm{d}V+\bar{g}_{ij}(U)\mathrm{d}Y^{i}\mathrm{d}Y^{j},$ (A.3)
where $\bar{g}_{ij}(U)=g_{ij}(U,0,0)=\lim_{\varepsilon\rightarrow 0}g_{ij}(U,\varepsilon^{2}V,\varepsilon Y^{k})$. This metric form is known as a plane wave metric in Rosen coordinates. It is worth noting that the finite region described by the cubic domain of the coordinates $(U,V,Y^{k})$ for the PL metric corresponds to the infinitely thin region along the geodesic $\gamma$ with a finite length of the affine parameter $U$. Therefore, taking the PL, we can extract the geometry in the vicinity of the null geodesic $\gamma$.
@@
$\mathrm{d}\bar{s}^{2}=2\mathrm{d}u\mathrm{d}v+A_{ab}(u)x^{a}x^{b}\mathrm{d}u^{2}+\delta_{ab}\mathrm{d}x^{a}\mathrm{d}x^{b},$ (A.4)
@@
$\bar{E}^{i}_{(b)}\partial_{u}\bar{E}^{(a)}_{i}=\bar{E}^{i}_{(a)}\partial_{u}\bar{E}^{(b)}_{i}$. (A.8)
Then the metric component $A_{ab}$ is given by
$A_{ab}=\delta_{ac}\ddot{\bar{E}}^{c}_{i}\bar{E}_{b}^{j}.$ (A.9)
@@
$g_{\mu\nu}E^{i}_{(a)}(\partial_{i})^{\nu}E^{j}_{(b)}(\partial_{j})^{\nu}=E_{j}^{(a)}E^{j}_{(b)}=\delta^{a}_{b},$ (A.10)
$(\partial_{U})^{\mu}\nabla_{\mu}(E^{i}_{(a)}(\partial_{i})^{\nu})=0.$ (A.11)
@@
The identification between $\bar{E}^{(a)}_{i}$ and $E^{(a)}_{i}$ enable us to make a covariant description of the PL. The covariant description can be made based on the following two facts:
$\bar{R}_{UiU}^{~~~~j}=\left.R_{UiU}^{~~~~j}\right|_{\gamma},$ (A.12)
$A_{ab}=-\bar{R}_{UiUj}\bar{E}^{i}_{(a)}\bar{E}^{j}_{(b)}=-\left.\left(R_{UiUj}E^{i}_{(a)}E^{j}_{(b)}\right)\right|_{\gamma},$ (A.13)
where $\bar{R}$ describes the Riemann tensor for $\bar{g}_{ij}$. We note that the first equation follows the fact that $R_{UiU}^{~~~~j}$ for the metric (A.1) does not depend on $a(U,V,Y^{k})$ nor $b_{i}(U,V,Y^{k})$. Since the PL metric in the Brinkmann coordinates is totally characterized by $A_{ab}$, we can obtain the metric form of Eq. (A.4) by calculating the Riemann tensor components and the parallelly propagated zweibein in the metric (A.1) without performing the explicit coordinate transformations.
@@ 2. Construction of the Penrose coordinates based on the Hamilton-Jacobi function
$g_{UU}$ $=$ $g_{\mu\nu}\frac{dx^{\mu}}{d\tau}\frac{dx^{\nu}}{d\tau}=0,$ (A.17)
$g_{UV}$ $=$ $g_{\mu\nu}\frac{\partial x^{\mu}}{\partial U}\frac{\partial x^{\mu}}{\partial V}=\partial_{\mu}S\frac{\partial x^{\mu}}{\partial V}=\frac{\partial S}{\partial V}=1,$ (A.18)
$g_{Ui}$ $=$ $g_{\mu\nu}\frac{\partial x^{\mu}}{\partial U}\frac{\partial x^{\mu}}{\partial Y^{i}}=\frac{\partial S}{\partial Y^{i}}=\frac{\partial V}{\partial Y^{i}}=0.$ (A.19)
@@ Appendix B Penrose limit for a circular photon orbit
$\mathcal{L}=\frac{1}{2}\left(-f\dot{t}^{2}+h\dot{r}^{2}+r^{2}\dot{\theta}^{2}\right).$ (B.1)
@@
$S(x^{\mu})=-Et+L\theta+\mathcal{\mathcal{}}R(r).$ (B.2)
In Ref. [17], the coordinate construction is given for the Schwarzschild metric with the initial hyper-surface being $r={\rm const.}$ surface. However, this surface $r={\rm const.}$ is tangent to the characteristic curve $r=r_{\rm c}$, and does not match the case we suppose. Although the procedure shown in Ref. [17] gives the same PL metric in Brinkmann coordinates as ours, we avoid this subtlety by taking a different initial hyper-surface. Specifically, we implicitly assume that the initial hypersurface is given by $F(\theta,r)=0$.
@@
$\mathrm{d}\theta$ $=$ $\dot{\theta}\mathrm{d}U+(\dot{\theta}-\beta)\mathrm{d}\tilde{r}$, (B.3)
$\mathrm{d}r$ $=$ $\dot{r}\mathrm{d}U+\dot{r}\mathrm{d}\tilde{r}$, (B.4)
@@
$\mathrm{d}S=\mathrm{d}V=-E\mathrm{d}t+L\mathrm{d}\theta+R^{\prime}(r)\mathrm{d}r$ (B.5)
$\Leftrightarrow$ $\mathrm{d}t=-\frac{1}{E}\mathrm{d}V+\frac{\beta^{2}\Delta^{2}}{E^{2}}\mathrm{d}U+\left[\frac{\Delta^{2}}{Efr}-\frac{L}{E}\left(\frac{L}{r^{2}}-\beta^{2}\right)\right]\mathrm{d}\tilde{r}$, (B.6)
@@
$\mathrm{d}\bar{s}^{2}$ $=$ $2\mathrm{d}U\mathrm{d}V+\frac{\beta^{2}\Delta^{2}}{E^{2}}\mathrm{d}\tilde{r}^{2}+r^{2}\sin^{2}\theta\mathrm{d}\tilde{\phi}^{2}$ (B.8)
@@
$\mathrm{d}\bar{s}^{2}=2\mathrm{d}U\mathrm{d}V+\frac{\beta^{2}\Delta^{2}(r(U))}{E^{2}}\mathrm{d}\tilde{r}^{2}+r^{2}(U)\sin^{2}\left(L\int r^{-2}(U)\mathrm{d}U\right)\mathrm{d}\tilde{\phi}^{2},$ (B.9)
@@
$E_{\mu}^{(\tilde{r})}=\frac{\beta\Delta(r(U))}{E}(\mathrm{d}\tilde{r})_{\mu}~~\&~~E_{\mu}^{(\tilde{\phi})}=r(U)\sin\left(L\int r^{-2}(U)\mathrm{d}U\right)(\mathrm{d}\tilde{\phi})_{\mu}.$ (B.10)
The metric components $A_{ab}$ can be easily calculated using Eq. (A.9). In the limit approaching a UCOP, we obtain
$A_{11}$ $=$ $\left.\frac{\ddot{\Delta}}{\Delta}\right|_{r=r_{\rm c}}=-V_{c}^{\prime\prime}=\alpha^{2},$ (B.11)
$A_{22}$ $=$ $\left.\frac{(r\sin\theta)\ddot{~}}{(r\sin\theta)}\right|_{r=r_{\rm c}}=-\left.\dot{\theta}^{2}\right|_{r=r_{\rm c}}=-\beta^{2}.$ (B.12)
@@
$t$ $\hat{=}$ $\frac{E}{f_{\rm c}}(u+\Lambda(v))-\frac{v}{E},$ (B.13)
$\theta$ $\hat{=}$ $\beta(u+\Lambda(v)),$ (B.14)
where we have used the symbol $\hat{=}$ to indicate the equality only on the UCOP. These relations correspond to Eqs. (2.19) and (2.22).